How do you solve the equation 3/2x=6

Found 2 solutions by chibisan, checkley77:
Answer by chibisan(131)

by cross multiply
3=6*2x
3=12x
1/4=x

Answer by checkley77(12569)

3/2x=6
3X/2=6
3X=6*2
3X=12
X=12/3
X=4 ANS.
PROOF:
3/2*4=6
12/2=6
6=6